OpenStudy (anonymous):

If a and b are the non-zero distinct roots of x^2 +ax +b=0, then the least value of the expression ax^2+bx+c is ??

OpenStudy (anonymous):

a > 0 right

OpenStudy (anonymous):

lol right or wrong @shubham.bagrecha

OpenStudy (anonymous):

why dont you post the solution

OpenStudy (anonymous):

@mukushla

OpenStudy (anonymous):

@ganeshie8 solving??

ganeshie8 (ganeshie8):

nope.. sorry just watching

OpenStudy (anonymous):

do u know the formulas for sum and product of roots for a quadratic equation?

OpenStudy (anonymous):

If a and b are non zero distinct roots then: \[D > 0\] \[b^2 - 4ac > 0\] Does that help??

OpenStudy (anonymous):

From this we get: \[a^2 > 4b\]

OpenStudy (jiteshmeghwal9):

i think that one of the roots of the given equation value is the smallest value.

OpenStudy (phi):

how about (x-a)(x-b)= x^2 -(a+b)x+ab must equal x^2+ax+b so match coefficients ab=b means a=1 -(b+1)= 1 b= -2 now we can find the min value of ax^2+bx+c

OpenStudy (anonymous):

@nitz

OpenStudy (anonymous):

i got it.